OpenStudy (bloomlocke367):

During your first week of training for a marathon, you run a total of 10 miles. You increase the distance you run each week by 20%. How many miles do you run during your twelfth week of training?

OpenStudy (bloomlocke367):

Would I use:\[a _{n}=a _{1}\times r ^{n-1}\]

OpenStudy (bloomlocke367):

and the common ratio is .2? because it's 20 PERCENT?

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (bloomlocke367):

@phi

OpenStudy (phi):

if you use 0.2 for r, the distance will get smaller for large n. you want to use 1.2 for r

OpenStudy (anonymous):

working on it

OpenStudy (bloomlocke367):

\[a _{n}=10\times 1.2^{n-1}\]

OpenStudy (bloomlocke367):

so that?

OpenStudy (phi):

you know the first week is 10 you know the second week is 20% more or 0.2*10= 2 more , so 10+2= 12

OpenStudy (anonymous):

20% of 10 is 2.0

OpenStudy (anonymous):

so ad that every week

OpenStudy (phi):

now test your equation, to make sure it "works" at n=1 (first week) you get 10 Good! at n=2 you get 10*1.2 = 12 also good. I would say that is a good equation

OpenStudy (phi):

how far you run is just your equation with n=12

OpenStudy (anonymous):

its 32 miles

OpenStudy (phi):

@qwertyboy56 the growth is exponential , so it will be larger than 32

OpenStudy (bloomlocke367):

so, it's\[a _{12}=10\times 1.2^{12-1}\] \[a _{12}=10\times 1.2^{11}\] \[=10\times 7.430083707\] \[=74.30083707\]

OpenStudy (bloomlocke367):

right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

i dont think its asking for a total

OpenStudy (phi):

you run 74 miles on the 12 week we could make a table (write down the distance each week, multiply by 1.2 to get the next week...) we will see it shoot up to 74 miles per week on the 12 week

OpenStudy (bloomlocke367):

Okay... Thank you so much! I just checked my answer in the back of the book... I just wanted to make sure I was doing it correctly, or I'd never learn XD